OpenStudy (debpriya):

There are 8 milks samples- A- 30%, B-70%,C-68%,D-59%,E-90%,F-97%,G-83%, H 37%. when two samples are mixed , how many mixtures will NEVER give youv63% milk.

Parth (parthkohli):

Simply when we have either of these cases: - Both samples are > 63% - Both samples are < 63% The only way to obtain 63% milk is when one exceeds that and one does not. Do you see what I mean?

Parth (parthkohli):

Samples > 63%: B, C, E, F, G Samples < 63%: A, D, H So your total count is 5C2 + 3C2 = 10 + 3 = 13.

OpenStudy (debpriya):

can you please explain with an example when can we obtain 633% of milk ?

OpenStudy (debpriya):

63% *

OpenStudy (debpriya):

suppose we take d and e where d is 59% and e is 90 % so If I mix these two how will I get 63 % ?

Parth (parthkohli):

Sorry - I have to go in a minute but We can get 63% milk when one of the samples >63% and the other <63% are mixed in some given ratio.

Parth (parthkohli):

For example, for 90 and 59, you can actually calculate the ratio. Let's say the ratio 59%:90% is \(k:1\). Then\[63 = \frac{59k + 90}{k+1}\]

Parth (parthkohli):

But obviously if you mix something like 90% and 68%, there is no way to get 63% milk. The purity will have to lie somewhere between 68% and 90%.

OpenStudy (debpriya):

oh okay ...thank you so much for your help :)

Parth (parthkohli):

No problem! =)